# Trends and Reversion in Financial Markets on Time Scales from Minutes to Decades

### v1

@@ 1 Introduction
Trendfollowing has been one of the most popular trading strategies in financial markets at least since the success of the turtle traders [1]. Market trends have also been studied in the literature since the 1990’s [2, 3, 4] and in many subsequent articles (see, e.g., [5, 6, 7, 8]). Traditional trend-followers are “long” a given market when its current trend is positive, and “short” when the trend is negative. The implicit assumption is that the future expected return in a given market has the same sign as the current trend in that market.
In [9], we have tested this assumption by a precise empirical measurement of the future expected return in a market as a function of its current trend strength, thereby confirming and extending previous observations in [10]. The key to filtering out the noise and getting high precision results was to aggregate not only across 30 years of daily futures returns and across a diversified set of equity indices, interest rates, currencies and commodities, but also across trend horizons ranging from a few days to a year. The latter aggregation makes sense, as it has long been known that similar trend-following strategies tend to work over this wide range of time scales.
The key results are shown in fig. 1: trends tend to revert before they become statistically strongly significant. In other words, by the time a trend has become so obvious that everybody can see it in a price chart, it is already over. This is consistent with the hypothesis that any obvious market inefficiency is quickly eliminated by investors. More precisely, for a given market and a given time horizon, we have defined the strength $\phi$ of a trend as its $t$-statistics. We have found that tomorrow’s return $R(t+1)$ (normalized to have variance 1) is modeled well by a cubic polynomial of today’s trend strength:
@@
where $a$ is a risk premium, $\epsilon$ represents random noise, and $b$ and $c$ are kinetic coefficients, which we have measured quite precisely. We have interpreted $b$ as the persistence of trends, and $c$, which is negative, as the strength of trend reversion, and observed the following:
  • The parameters $b,c$ are universal in the sense that they seem to be the same for all assets within the limits of statstical significance. This is in line with the fact that many successful trend-followers use the same systematic trading strategy for all assets.
  • $c$ is negative and does not seem to depend much on the trend’s time horizon $T$. However, $b$ depends on the horizon: it peaks at $T\sim$ 3-12 months, which is in line with the time scales on which trendfollowers typically operate. $b$ decays for longer or shorter horizons and appears to become negative for $T<1$ day or $T>$ several years.
@@
Figure 1: Left: The expectation value $E(r)$ of the next day’s return of a futures market can be modeled by a cubic polynomial of the $t$-statistics $\phi$ of the trend. The linear term $b\phi$ models trend-persistence. The cubic term $c\phi^{3}$ with $c<0$ models trend-reversion. Right: The coefficients $b$ and $c$ depend on the time horizon $T$, which runs from 2 days to 4 years.
Studying the interplay of trends and reversion in financial markets is not only of interest for traders. It also holds clues for developing an accurate model of the markets that can replicate these and other empirically observed properties, such as the so-called “stylized facts” of finance [11, 12]. Such a comprehensive financial market model is in turn needed for better risk management, including the monitoring and managing of systemic risk.
Indeed, the above empirical observations have led us to propose a lattice gas model of financial markets [13]. The lattice represents the social network of investors, while the gas molecules represent the shares of an asset. The order parameter, i.e., the deviation of the molecule density from its mean, is identified with the deviation of the asset price from its intrinsic value. Market efficiency implies a form of self-organized criticality [14]: one expects arbitrageurs to drive this lattice gas to its critical point, where a second-order phase transistion occurs. Near this point, the model can indeed explain the stochastic process (1), where the cubic driving force is the derivative of the so-called Landau potential known from the theory of phase transitions. This model can also explain other observed analogies between financial markets and critical phenomena, including the universality of the coefficients $b,c$.
The quantitative predictions of such a network model of financial markets depend on the precise properties of the network, such as its topology or its fractal dimension. Different network models predict different scaling behaviors of the markets under a rescaling of the time horizon $T$. For example, if a simple (and unrealistic) hypercubic network topology was assumed, then a fractal dimension of $D\approx 3$ could quantitatively explain the observed value of the second Hurst coefficient of market returns [13]. The model predictions for more realistic network topologies, such as in [15], are being worked out seperately.
In the current paper, we focus on establishing the empirical facts, i.e., the observed scaling bahaviour that such theoretical models must reproduce: we extend the empirical measurement of the parameters $b,c$, as plotted in fig. 1 (right), in two directions:
@@
  • to long-term scales from years to decades, using financial data since medieval times, extending previous historical results [16, 17, 18] to trends lasting longer than a year.
This note is organized as follows. In section 2, we briefly review the methodology already applied in [9] to daily data. In section 3, we extend these results to intraday data. In section 4, we extend them to long-term data. Section 5 combines and summarizes our observations and interprets them within the framework of the lattice gas model [13].
2 Review and Methodology
To empirically measure the interplay of trends and reversion in financial markets, we follow the methodology of [9, 19], which we briefly review in this section.
2.1 Data
The analysis of daily and intra-day market data is based on historical log-returns of a set of futures prices. We use futures prices, instead of the underlying market prices, because futures prices are guaranteed to be marked-to-market. Our set of futures is diversified across asset classes (equity indices, interest rates, currencies, commodities), regions (Americas, Europe, Asia) and commodity sectors (energy, metals, agriculture).
For daily closing prices, we have used the broadly diversified set of liquid contracts in table 1 over the period 1990-2020 [19]. For intraday data, we will use end-of-minute prices of a similar set of futures contracts described in section 3. For long-term data covering several centuries, we will use end-of-month prices of a diversified set of assets as described in section 4, as the histories of futures prices are insufficient. For an indicative analysis of very lomg-term time horizons, we also use yearly data since medieval times for some assets.
Table 1: Markets America Europe Asia
@@ 2.2 Trend Definition
For each market $i$ and each time intervall, we consider the closing prices $P_{i}(t)$. We define normalized log-returns $R_{i}(t)$:
@@
where the long-term risk premium $\mu_{i}$ and the long-term standard deviation $\sigma_{i}$ of a market $i$ are measured over the whole time period (to avoid biases, in the out-of-sample cross-validation of our results, the risk premia and variances are estimated only from the training samples, excluding the validation samples). For each market and each time interval, we compute different trend strengths on different (partly overlapping) time scales $T$:
@@
At a given point in time, a market may well be in an up-trend on one scale, but on a down-trend on another scale. For a given horizon $T$, we define the trend strength $\phi_{i,T}(t)$ of a given market $i$ at the close of trading of interval $t\in Z$ as a weighted average of previous log-returns of that market in excess of the long-term risk premium:
$\phi_{i,T}(t)=\sum_{n=0}^{\infty}w_{T}(n)\cdot\hat{R}_{i}(t-n)\ \ \ \text{with}\ \ \ \hat{R}_{i}(t-n)=R_{i}(t-n)-{\mu_{i}\over\sigma_{i}},$
where $w_{T}(n)$ is a weight function. For monthly data, we use a variant of these trend strengths, where the volatilities are computed by an exponentially weighted moving average over the preceding $30$ years. We normalize the weight function $w_{T}(n)$ such that the trend strength $\phi_{i,T}$ has standard deviation 1. Assuming that returns over non-intersecting intervals are independent of each other (which is true to high accuracy), this implies:
$\sum_{n=0}^{\infty}w_{T}^{2}(n)=1.$
@@
$\tilde{w}_{T}(n)$ $=$ $M_{T}\cdot\ n\cdot e^{-2n/T}\ \ \ \text{with normalization factor}\ \ M_{T}.$
@@
Table 2 shows a small excerpt of the resulting data table in the case of $N$ daily returns. It contains $N\times 24\times 10$ pairs of trend stregths on a given day and the next-day return.
Table 2 Trend strength at given day Return
@@ 2.3 First Look at the Data
For a first explorative data analysis, we define buckets of similar trend strengths and measure the average future return within each bucket. For daily data, fig. 1 (left) shows the average next-day return as a function of the trend strength on a given day. The results are aggregated across 30 years of daily returns for each market, and across all 24 markets. Unlike previous studies, [9] also aggregates across different time scales $T$. This was the key to clearly observe the statistically highly significant patterns in fig. 1. For shorter time windows, or for any single time scale, these patterns are too blurred to be detected reliably. This may explain why our findings have not been reported already a decade ago.
@@ 2.4 Regression Analysis
To prove and quantify this suspected pattern, we dissolve our trend strength buckets again and perform a polynomial regression on the full underlying data set of $24\cdot 260\cdot 30\cdot 10\approx 2$ million data pairs (24 markets, 260 trading days per year, 30 years, 10 trend strengths). We regress the next-day normalized log-return $R_{i}(t+1)$ (2) against a polynomial of the current trend strength $\phi_{i,k}(t)$ across all markets $i$ and time scales $T=2^{k}$.
@@
where $\epsilon$ represents random noise, and $a$ measures the average normalized risk premium $\mu_{i}/\sigma_{i}$ across all assets. We ignore $a$ and focus on determining the persistence of trends $b$, and the strength of trend reversion $c$. The results of the overall regression analysis are shown in table 3. Quadratic and higher-order terms in $\phi$ are found not to be statistically significant.
Table 3 Daily regression with linear and cubic terms
@@ 2.5 Statistical Analysis
  • The error of the coefficients and their t-statistics were computed by bootstrapping, using 5000 bootstrapping samples. Regression on each new sample of days yielded the distribution of the 5000 regression coefficients $b$ and $c$ shown in fig. 2 (right), from which the estimation errors in table 3 were read off. These actual errors are 3-4 times as large as the errors that standard regression tools would report. Nevertheless, even with these larege errors the regression coefficients are still statistically highly significant.
  • The adjusted R-squared was computed by 15-fold cross validation. The resulting out-of-sample R-squared is reported as $R^{2}_{adj}$ in table 3. The actual out-of-sample correction $R^{2}-R^{2}_{adj}$ is about 3 times as big as what standard regression tools would report. Nevertheless, the out-of-sample $R^{2}_{adj}$ is still highly significant despite being very small, and agrees with the Sharpe ratio that has been achieved by trend-followers in practise.
@@
The regression results of table 3 and further regression analyses in [9] confirm and quantify our conclusions from fig. 1. In particular, we see that the values of $b$ and $c$ although very small - are statistically highly significant, despite the fact that market returns are neither normally, nor identically, nor independently distributed. The reported values of $b$ and $c$ are averaged over the past 30 years. In more recent time periods, our measurements indicate that at least $b$ has decreased. This has a natural interpretation: so much money has flown into trend-following that it does not work as well any more as it used to.
3 Intraday Time Scales
We now extend the previous analysis from daily data to intraday data with time scales ranging from 2 minutes to 1 trading day, using tick data from world-wide futures markets. For previous work on intraday trends see, e.g., [20, 21].
3.1 Data
As in [19], our analysis is based on historical log-returns of futures contracts that are rolled 5 days prior to first notice. We again diversify across four different asset classes (equity indices, interest rates, currencies, and commodities) and three regions (Americas, Europe, and Asia) as shown in table 4. Compared with the analysis of daily data, we drop some markets that are less liquid or for which we have less clean data (Nikkei and Hang Seng indices, Japanese and Australian government bonds, AUD and NZD currencies, copper and live cattle). We also replace the DAX by the Eurostoxx 50 Index. Our data were provided by TickData (https://www.tickdata.com) and cover the 14-year period 2010-2023.
Table 4 Market 1 Market 2 Market 3 Market 4
@@ 3.1 Data
The analysis is based on end-of-minute prices (last traded price in each minute), rather than the whole set of tick data which records every individual transaction with a time stamp These end-of-minute prices are sufficient for our purposes, as we do not analyze intra-minute time scales here. From the prices, we compute the returns following the definition (2), normalizing them in the same fashion as in [19] to have standard deviation 1.
The analysis of intraday data must account for the overnight jump in market prices due to closure of the exchange and re-opening on the next trading day. As a result, the return in the opening minute is typically quite large, which can distort our analysis. Moreover, in the absence of overnight data, no meaningful trend strength can be computed during the night and in the early trading day. We therefore analyze the intraday trending behavior for trend horizons of $2^{k}$ minutes with $k\in\{1,2,...,6\}$ ($k=6$ corresponding to roughly 1 hour) on a day-by-day scheme: the time series of returns are broken up into sub-sets, one for each day, from the opening to the close of each market in its respective time zone. On each day, the first-minute return is ignored, and the first $T$ minutes of the day serve to build up the $T$-minute trend strength. After this ramp-up period, the trend strengths of the different horizons at the end of each minute as well as the next-minute return are measured.
@@
This day-by-day analysis limits the trend horizons $T=2^{k}$ minutes to at most a few hours. For $k\geq 7$, we instead use continuous time series of minute returns through the whole 14-year period, reinstating the first-minute returns to include the overnight returns. When connecting minute results with daily results, we interpret $2^{10}$ minutes as one trading day.

### v2

@@ 1 Introduction
Trendfollowing has been one of the most popular trading strategies in financial markets at least since the success of the turtle traders in the 1980’s [1]. Market trends have also been studied in the literature since the 1990’s. Early studies such as [2, 3] have shown that trend-following strategies can consistently generate risk-adjusted returns that exceed ordinary risk premia. Alternative beta strategies have tried to systematically replicate them [4, 5], and many analyses, such as [6, 7, 8, 9, 10, 11], have demonstrated that these excess returns arise across all asset classes, have low correlation with traditional asset returns, and complement other systematic strategies well. More recently, it has also been shown that market trends have already existed at least since the 19th century [12, 13, 14].
Traditional trend-followers are “long” a given market when its current trend is positive, and “short” when the trend is negative. The implicit assumption is that the future expected return in a given market has the same sign as the current trend in that market. This contrasts with the intuitive expectation that a trend should revert once it has taken the market price of an asset too far away from its intrinsic value. This trend reversion has been quantitatively observed in [13] and has been modelled in [15].
As reported in [16], we have independently confirmed and refined this observation by a precise empirical measurement of the future expected return in a market as a function of its current trend strength. The key to filtering out the noise and getting high precision results was to aggregate not only across 30 years of daily futures returns and across a diversified set of 24 equity indices, interest rates, currencies and commodities, but also across 10 trend horizons ranging from a few days to a year, as will be reviewed in more detail section 2. The latter aggregation makes sense, as it has long been known that similar trend-following strategies tend to work over this wide range of time scales.
Figure 1: Left: The expectation value $E(r)$ of the next day’s return of a futures market can be modeled by a cubic polynomial of the $t$-statistics $\phi$ of the trend. The linear term $b\phi$ models trend-persistence. The cubic term $c\phi^{3}$ with $c<0$ models trend-reversion. Right: The coefficients $b$ and $c$ depend on the time horizon $T$, which runs from 2 days to 4 years. Both figures are aggregated across 24 assets and 10 time horizons as discussed in the text.
A key result of this previous study [16], aggregated across assets and trend horizons, is shown in fig. 1 (left): trends tend to revert before they become statistically strongly significant. In other words, by the time a trend has become so obvious that everybody can see it in a price chart, it is already over. This is consistent with the hypothesis that any obvious market inefficiency is quickly eliminated by investors. More precisely, for a given market and a given time horizon, we have defined the strength $\phi$ of a trend as its $t$-statistics. We have found that tomorrow’s return $R(t+1)$ (normalized to have variance 1) is modeled well by a cubic polynomial of today’s trend strength:
@@
where $a$ is a risk premium, $\epsilon$ represents random noise, and $b$ and $c$ are kinetic coefficients. A similar model had already been proposed in [15] and extended in [17]. We have interpreted $b$ as the persistence of trends, and $c$, which is negative, as the strength of trend reversion, measured them quite precisely, and refined our observations as follows:
  • The parameters $b,c$ are universal in the sense that they seem to be the same for all assets within the limits of statistical significance. This is in line with the fact that many successful trend-followers use the same systematic trading strategy for all assets.
  • $c$ is negative and does not seem to depend much on the trend’s time horizon $T$. However, $b$ depends on $T$: it peaks at $T\sim$ 3-12 months, which is in line with the time scales on which trendfollowers typically operate. $b$ decays for longer or shorter horizons and may become negative for $T<1$ day or $T>$ several years (fig. 1, right).
@@
Studying the interplay of trends and reversion in financial markets is not only of interest for traders. It also holds clues for developing an accurate model of the markets that can replicate these and other empirically observed properties, such as the so-called “stylized facts” of finance [18, 19]. Such a comprehensive financial market model is in turn needed for better risk management, including monitoring and managing systemic risk.
Indeed, the above empirical observations have led us to propose a lattice gas model of financial markets [20]. The lattice represents the social network of investors, while the gas molecules represent the shares of an asset. The order parameter, i.e., the deviation of the molecule density from its mean, is identified with the deviation of the asset price from its intrinsic value. Market efficiency implies a form of self-organized criticality [21]: one expects arbitrageurs to drive this lattice gas to its critical point, where a second-order phase transition occurs. Near this point, the model can indeed explain the stochastic process (1), where the cubic driving force is the derivative of the so-called Landau potential known from the theory of phase transitions. This model can also explain other observed analogies between financial markets and critical phenomena, including the universality of the coefficients $b,c$.
The quantitative predictions of such a network model of financial markets depend on the precise properties of the network, such as its topology or its fractal dimension. Different network models predict different scaling behaviors of the markets under a rescaling of the time horizon $T$. For example, if a simple (and unrealistic) hypercubic network topology was assumed, then a fractal dimension of $D\approx 3$ could quantitatively explain the observed value of the second Hurst coefficient of market returns [20]. The model predictions for more realistic network topologies, such as in [22], are being worked out separately.
In the current paper, we focus on establishing the empirical facts, i.e., the observed scaling bahaviour that such theoretical models must reproduce: we extend the previous empirical measurement of the parameters $b,c$, as plotted in fig. 1 (right), in two directions:
@@
  • to long-term scales from years to decades, using financial data since medieval times, thereby extending previous historical results [12, 13, 14, 15].
This paper is organized as follows. In section 2, we briefly review the methodology already applied in [16] to daily data. In section 3, we extend these results to intraday data. In section 4, we extend them to long-term monthly data. Section 5 combines and summarizes our observations and interprets them within the framework of the lattice gas model of [20].
2 Review: Analysis of Daily Data
To empirically measure the interplay of trends and reversion in financial markets, we follow the methodology of [16] based on the data [23], which we briefly review in this section.
2.1 Data Set
The analysis of daily and intra-day market data is based on historical log-returns of a set of futures prices. We use futures prices, instead of the underlying market prices, because futures prices are guaranteed to be marked-to-market every day. Our set of futures is diversified across asset classes (equity indices, interest rates, currencies, commodities), regions (Americas, Europe, Asia) and commodity sectors (energy, metals, agriculture).
For daily closing prices, we have used the broadly diversified set of liquid contracts in table 1 over the period 1990-2020 [23]. For intraday data, we will use end-of-minute prices of a similar set of futures contracts, as described in section 3. For long-term data covering several centuries, we will use end-of-month prices of a diversified set of assets, as described in section 4, since the histories of futures prices are insufficient. In addition, for an indicative analysis of very long-term time horizons, we will use yearly data since medieval times for some assets in section 4.
Asset classes vs. regions America Europe Asia
@@ 2.1 Data Set
Table 1: List of financial markets used in the analysis of daily data.
@@ 2.2 Trend Definition
For each market $i$ and each time interval, we consider the closing prices $P_{i}(t)$. We define normalized log-returns $R_{i}(t)$:
@@
where the long-term risk premium $\mu_{i}$ (used below) and the long-term standard deviation $\sigma_{i}$ of a market $i$ are measured over the whole time period (to avoid biases, in the out-of-sample cross-validation of our results, the risk premia and variances are estimated only from the training samples, excluding the validation samples). For each market and each time interval, we compute different trend strengths on different (partly overlapping) time scales $T$:
@@
At a given point in time, a market may well be in an up-trend on one scale, but in a down-trend on another scale. For a given horizon $T$, we define the trend strength $\phi_{i,T}(t)$ of a given market $i$ at the close of trading of interval $t\in Z$ as a weighted average of previous log-returns of that market in excess of the long-term risk premium:
$\phi_{i,T}(t)=\sum_{n=0}^{\infty}w_{T}(n)\cdot\hat{R}_{i}(t-n)\ \ \ \text{with}\ \ \ \hat{R}_{i}(t-n)=R_{i}(t-n)-{\mu_{i}\over\sigma_{i}},$ (3)
where $w_{T}(n)$ is a weight function. For monthly data, we use a variant of these trend strengths, where the variances are computed by an exponentially weighted moving average over the preceding $30$ years. We normalize the weight function $w_{T}(n)$ such that the trend strength $\phi_{i,T}$ has standard deviation 1. Assuming that returns over non-intersecting intervals are independent of each other (which is true to high accuracy), this implies:
$\sum_{n=0}^{\infty}w_{T}^{2}(n)=1.$ (4)
@@
$\tilde{w}_{T}(n)$ $=$ $M_{T}\cdot\ n\cdot e^{-2n/T}\ \ \ \text{with normalization factor}\ \ M_{T}.$ (5)
@@
Table 2 shows a small excerpt of the resulting data table. It contains $N\times 24\times 10$ pairs of trend stregths on a given day, where $N$ is the number of days, 24 is the number of assets, and 10 is the number of trend horizons. The response variable is the next-day return.
Time / Market Trend strength at given day Return
@@
Table 2: Small excerpt of the data table used in the daily data analysis, showing two trading days (out of 30 years of data), and 3 of the 24 assets. For each asset and each day, 10 different trend strengths are computed, measuring the trend over the past $T=2^{k}$ days with $k\in\{1,2,...,10\}$. The last column is the response variable, the asset return on the day after.
@@ 2.3 First Look at the Data
For a first explorative data analysis, we define 15 buckets $b_{k}$ of similar trend strength
$\phi\in b_{k}\ \ \ \text{with}\ \ \ k\in\{-7,+7\},\ \ \ b_{k}=[\ {k\over 3}-{1\over 6}\ ,\ {k\over 3}+{1\over 6}\ [\ .$ (6)
[…]
@@ 2.4 Regression Analysis
To prove and quantify this suspected pattern, we dissolve our trend strength buckets (6) again and perform a polynomial regression on the full underlying data set of $24\cdot 260\cdot 30\cdot 10\approx 2$ million data pairs (24 markets, 260 trading days per year, 30 years, 10 trend strengths). The following results are thus independent of the definition of buckets in (6). We regress the next-day normalized log-return $R_{i}(t+1)$ (2) against a polynomial of the current trend strength $\phi_{i,k}(t)$ across all markets $i$ and time scales $T=2^{k}$.
@@
where $\epsilon$ represents random noise, and $a$ measures the average normalized risk premium $\mu_{i}/\sigma_{i}$ across all assets. We ignore $a$ and focus on determining the persistence of trends $b$, and the strength of trend reversion $c$. The results of the overall regression analysis are shown in table 3. As reported in the appendix, the $\phi^{2}$ term is not statistically significant, and adding higher-order terms in $\phi$ would also reduce the statistical significance of the model.
Parameter Nonlinear regression results
@@
Table 3: Regression results for daily returns: $a$ is the intercept, while $b$ and $c$ are the regression coefficients of the two explanatory variables, the trend strength and its cube.
@@ 2.5 Statistical Analysis
  • The error of the coefficients and their t-statistics were computed by bootstrapping, using 5000 bootstrapping samples. Regression on each new sample of days yielded the distribution of the 5000 regression coefficients $b$ and $c$ shown in fig. 2 (right), from which the estimation errors in table 3 were read off. These actual errors are 3-4 times as large as the errors that standard regression tools would report. Nevertheless, even with these large errors the regression coefficients are still statistically highly significant.
  • The adjusted R-squared was computed by 15-fold cross validation. The resulting out-of-sample R-squared is reported as $R^{2}_{adj}$ in table 3. The actual out-of-sample correction $R^{2}-R^{2}_{adj}$ is about 3 times as big as what standard regression tools would report. Nevertheless, the out-of-sample $R^{2}_{adj}$ is still highly significant despite being very small, and agrees with the Sharpe ratio that has been achieved by trend-followers in practice.
@@
The regression results of table 3 and further regression analyses in [16] confirm and quantify our conclusions from fig. 1. In particular, we see that the values of $b$ and $c$ although very small - are statistically highly significant, despite the fact that market returns are neither normally, nor identically, nor independently distributed.
The reported values of $b,c$ are averaged over the past 30 years. In more recent time periods, our measurements indicate that at least $b$ has decreased. E.g., section 4.4 of [16] confirms this at the 97.5% confidence level, based on a linear decay ansatz, and estimates that $b\rightarrow 0$ somewhere between 2010 and 2030. This has a natural interpretation: so much money has flown into trend-following that it does not work as well any more as it used to.
3 Analysis of Intraday Data
We now extend the previous analysis from daily data to intraday data with time scales ranging from 2 minutes to 1 trading day, using tick data from world-wide futures markets (i.e., data that record every buy order, every sell order, and every trade, including a precise timestamp). For previous work on intraday trends see, e.g., [24, 25].
3.1 Data Set
As in [23], our analysis is based on historical log-returns of futures contracts that are rolled 5 days prior to first notice (i.e., the expiring contract is sold and a new contract is bought). We again diversify across four different asset classes (equity indices, interest rates, currencies, and commodities) and three regions (Americas, Europe, and Asia) as shown in table 4. Compared with the analysis of daily data, we drop some markets that are less liquid or for which we have less clean data (Nikkei and Hang Seng indices, Japanese and Australian government bonds, AUD and NZD currencies, copper and live cattle). We also replace the DAX by the Eurostoxx 50 Index. Our data were provided by TickData (https://www.tickdata.com) and cover the 14-year period 2010-2023.
Asset Class Market 1 Market 2 Market 3 Market 4
@@ 3.1 Data Set
Table 4: List of financial markets used in the analysis of intraday (end-of-minute) data.
The analysis is based on end-of-minute prices (last traded price in each minute), rather than the whole set of intra-minute tick data. These end-of-minute prices are sufficient for our purposes, as we do not analyze intra-minute time scales here. From the prices, we compute the returns following the definition (2), normalizing them in the same fashion as in [23] to have standard deviation 1.
The analysis of intraday data must account for the overnight jump in market prices due to the closure of the exchange and re-opening on the next trading day. As a result, the return in the opening minute is typically quite large, which can distort our analysis. Moreover, in the absence of overnight data, no meaningful trend strength can be computed during the night and in the early trading day. We therefore analyze the intraday trending behavior for trend horizons of $2^{k}$ minutes with $k\in\{1,2,...,6\}$ ($k=6$ corresponding to roughly 1 hour) on a day-by-day scheme: the time series of returns are broken up into sub-sets, one for each day, from the opening to the close of each market in its respective time zone. On each day, the first-minute return is ignored, and the first $T$ minutes of the day serve to build up the $T$-minute trend strength. After this ramp-up period, the trend strengths of the different horizons at the end of each minute as well as the next-minute return are measured.
@@
This day-by-day analysis limits the trend horizons $T=2^{k}$ minutes to at most a few hours. For $k\geq 7$, we instead use continuous time series of minute returns through the whole 14-year period, reinstating the first-minute returns to include the overnight returns. When connecting minute results with daily results, we interpret $2^{10}$ minutes as one trading day. This accounts for the fact that there is no or little overnight trading in many markets.